The supply and demand equations for the production of phone cases are given below. The variables p and q represent price (in dollars) and quantity, respectively. Determine the quantity of phone cases when the market is in equilibrium.

The supply and demand equations for the production of phone cases are given below. The variables p and q represent price (in dollars) and quantity, respectively. Determine the quantity of phone cases when the market is in equilibrium.

Supply: p = −2q + 180
Demand: p = −4q − 120

150

10

30

50

None of these
Answer #1

Given Supply : p =-22 +180 Given Demand: p=-42 - 120 market is in equilibrium then demand is equal to supply and therepre qua